OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

your welcome

OpenStudy (arabpride):

I think it's better if you explain how you got that answer .... this is a site to help you learn... not for giving answers.... please read the CoC

OpenStudy (arabpride):

-3x - 8 + 6x = 7 combine like terms: -3x - 8 + 6x = 7 -3x & 6x 3x - 8 = 7 +8 +8 3x = 15 /3 /3 x = 5 。◕‿◕。

OpenStudy (anonymous):

Thank you

OpenStudy (arabpride):

Welcome *^.^*